OpenStudy (atomicreaper):

Consider the equation 3tan^2(x)-1=0.

Nnesha (nnesha):

solve for x ? find exact solutions ??

OpenStudy (dayakar):

60degrees

Nnesha (nnesha):

oh yea alright solve for tan

OpenStudy (atomicreaper):

Im still trying to figure it out really.

Nnesha (nnesha):

ah i see 2pi/3 5pi/3 is ur answer ??

OpenStudy (dayakar):

\[3\tan ^{2}x=1\]

OpenStudy (dayakar):

\[\tan ^{2}x=1/3\]

OpenStudy (atomicreaper):

No its just hat i had accidentally clicked

Nnesha (nnesha):

what's the rush ?

OpenStudy (atomicreaper):

Like I need it in the form that is on the questions, I really almost have no Idea what im doing here.

Nnesha (nnesha):

alright first we have to solve for tanx someone already solved first 2 steps

Nnesha (nnesha):

\[\huge\rm 3\tan^2(x)-1=0\] to solve for tan first add one both sides we need tan itself at left side so divide both sides by 3 \[\tan^2(x) =\frac{1}{3}\]

Nnesha (nnesha):

what would be the next step ? how would you cancel out the square from tan^2 ?

OpenStudy (atomicreaper):

Would you separate it where it is: (tan)(tan) and cancel one making it just Tan? Like I said I'm kind of lost, so I'm not entirely sure.

Nnesha (nnesha):

let tan(x) =y \[\huge\rm y^2=\frac{1}{3}\] do you know wha would you do to cancel out the square ??? :=))

Nnesha (nnesha):

let me know if u r not sure :=))

OpenStudy (atomicreaper):

I am not entirely sure, I am way better at english. . . I hate math.

Nnesha (nnesha):

hm alright to cancel out the square we should take `square root` square and square cancelz each other out \[\sqrt{x^2} = x\]

Nnesha (nnesha):

so take square root `both sides ` of the equal sign

OpenStudy (atomicreaper):

ok, so it would be \[√x^2=x^2\]?

OpenStudy (atomicreaper):

Wait

OpenStudy (atomicreaper):

\[√x^2=√x\]

Nnesha (nnesha):

hmm no square and square cancels each other ot  this is an example btw